The sum of magnitude of two force acting at a point is 16 N.If the resultant force is 8N and its direction is perpendicular to smaller force then The forces are

Grade:12th pass

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
Let the 2 forces be P and Q and R is resulatant force perpendicular to P.
so given, P+Q=16 (P
from PQthagoras' Theorem, R^2=Q^2-P^2
=>R^2=(P+Q) (Q-P) =16(Q-P)
=>(Q-P)=R^2/(Q+P)=8^2 /16 =4
solving Q+P=16 and Q-P=4, we get
P=6 N and Q=10 N [Ans]
Regards,
Thank You,
Soumya Ranjan Mohanty.